OpenStudy (anonymous):

A peregrine falcon dives at a pigeon. The falcon starts downward from rest with free-fall acceleration. If the pigeon is 56.0 m below the initial position of the falcon, how long does the falcon take to reach the pigeon? Assume that the pigeon remains at rest. ______S

OpenStudy (mathmate):

Poor pigeon, she doesn't know what's ahead. \[S\ =\ \frac{1}{2} g t^2\] S= vertical distance g=acceleration due to gravity = 9.81 m/s^2 t=time in seconds. Solve for t.

OpenStudy (anonymous):

can u use draw and plug in the equation for me

OpenStudy (anonymous):

?

OpenStudy (mathmate):

 S=56m t=time (to be determined) g=9.81 m/s^2 56=(1/2)9.81 t^2 t^2=2*56/9.81 can you find t (time)?

OpenStudy (mathmate):

@btwosick is everything ok?